# John Mendonça1,2, Isabel Trancoso1,2    Alon Lavie3,4

### v3

@@ C.3 Statistics
C.4 Translation vs Native Generation
@@ D.2 Additional analysis
Table 3: Inter-Annotator Agreement (IAA) for all dimensions across languages with 2 annotators. A formal definition of these dimensions is available in Table 1.
@@ E.1 Human Annotations
We present Inter-Annotator Agreement (IAA) agreement scores in Table 3, calculated with using Krippendorff’s $\alpha$ [^13] We omit bootstrap intervals due to instability caused by the low number of positive examples in an overall small sample size (100), which leads to unreliable or undefined CI.. Issues such as non-factual (!Fac) and unsafe (!Safe) achieve very high correlations. This is due to the relatively objective nature of this issue [^14] Furthermore, annotators were allowed to consult external references to check for ”factuality”.. Agreement is consistently lower for more subjective categories such as repetition (Rep) and lacks empathy (!Emp), highlighting the challenges of reliably annotating these issues.
@@ E.3 Additional Examples
We present additional examples from our curated meta-evaluation benchmark in Figures 14, 15, 16 and 17.

### v4

@@ C.4 Statistics
C.5 Translation vs Native Generation
@@ D.2 Additional analysis
EN 1.00 0.95 0.91 0.97 0.76 0.82 0.79 0.94 0.84
ZH 0.76 0.98 0.94 0.96 0.85 0.74 0.77 0.96 0.68
FR 0.94 0.96 0.97 0.98 0.97 0.95 0.92 0.96 0.76
PT 0.82 0.93 0.90 0.91 0.74 0.68 0.82 0.95 0.82
Table 3: Inter-Annotator Agreement (IAA) calculated using the percentage of exact agreement between annotators across all dimensions and languages with two annotators. For "overall", we calculate the adjacent agreement rate, i.e., the percentage of score pairs that differ by no more than 1 (liu2025mdsevalmetaevaluationbenchmarkmultimodal). A formal definition of these dimensions is available in Table 1.
Language !Int !Safe !Emp !Com Rep !Coh !Rel !Fac Overall
@@
Table 4: Inter-Annotator Agreement (IAA) calculated using Krippendorff’s $\alpha$ for all dimensions across languages with two annotators.
@@ E.1 Human Annotations
We present agreement rates across all dimensions and languages for which we had more than one annotator in Table 3. Following Finch et al. (2023a), we additionally present IAA calculated using Krippendorff’s $\alpha$ [^13] We omit bootstrap intervals due to instability caused by the low number of positive examples in an overall small sample size (100), which leads to unreliable or undefined CI. We additionally report raw percentage agreement (Table 3). in Table 4. Agreement values are generally within the agreement reported by Finch et al. (2023a) for behaviour labels ($0.3\leq\alpha\leq 0.8$)
Overall, issues such as non-factual (!Fac) and unsafe (!Safe) achieve very high correlations. This is due to the relatively objective nature of this issue [^14] Furthermore, annotators were allowed to consult external references to check for ”factuality”.. Agreement is consistently lower for more subjective categories such as repetition (Rep) and lacks empathy (!Emp), highlighting the challenges of reliably annotating these issues.
@@ E.3 Additional Examples
We present additional examples from our curated meta-evaluation benchmark in Figures 15, 16, 17 and 18.